OpenStudy (artfis):

4x -9 = 4x - 7 Is there even a solution?

hartnn (hartnn):

4x -9 = 4x - 7 gives you -9=-7 which is NEVER true so, there is no solution to 4x -9 = 4x - 7

OpenStudy (artfis):

Ok......now I have both sides in Absolute Value. Does that change it?

hartnn (hartnn):

yes.

hartnn (hartnn):

because you can have (4x-9) = - (4x-7)

hartnn (hartnn):

when |x|=a x=a, x=-a

hartnn (hartnn):

(4x-9) = - (4x-7) <-----has a solution can you find it ?

OpenStudy (artfis):

x = 2

OpenStudy (artfis):

I'm confused if it is no solution when it is 4x - 9 = 4x - 7 though

hartnn (hartnn):

x=2 is correct :) wats the confusion ? -9 can never equal -7 so, 4x-9=4x-7 cannot hold true for any value of x.

OpenStudy (artfis):

Oh ok. Thank you I see my problem. I just needed confirmation. Thank you for your help :D

hartnn (hartnn):

welcome ^_^